OpenStudy (anonymous):

A rock is dropped from a sea cliff and the sound of it striking the ocean is heard 4.2 s later. If the speed of sound is 340 m/s, how high is the cliff?

OpenStudy (amistre64):

is it heard 4.2 s after releasing the rock, or after seeing it hit the water?

OpenStudy (anonymous):

it is heard 4.2 s after releasing the rock

OpenStudy (amistre64):

hmm, then we would need to develop a gravity model i spose

OpenStudy (anonymous):

all I know is g=9.80 m/s^2

OpenStudy (anonymous):

I switched into the class a week late so i am behind on these problems

OpenStudy (amistre64):

hmm, h(t) = -4.9 t^2 + h would be the falling equation

OpenStudy (anonymous):

how did you get -4.9?

OpenStudy (amistre64):

9.8/2

OpenStudy (amistre64):

do you know how to work calculus?

OpenStudy (anonymous):

i am taking calculus this year

OpenStudy (amistre64):

well, what you may learn is: acceleration, integrates to velocity, integrates to distance as such, if a = k v = kt + c s = kt^2/2 +ct + d since we know k=-9.8 (its a downward force), and the initial velocity is 0, so c=0. the d is the only unknown

OpenStudy (amistre64):

now, we know the total time it takes is 4.2 seconds ..... and the height going down is the same as it coming up then we can start to equate things and hopefully it all works out

OpenStudy (anonymous):

so plug 4.2 into t?

OpenStudy (amistre64):

lets try: d = -4.9t^2 and d = 340t the time it takes to fall is not the same time it takes for the sound to get there so we cant just plug in 4.2

OpenStudy (amistre64):

im thinking the negative is not going to be proper here. so lets try: 4.9t^2 = 340t 4.9t^2 - 340t = 0 t(4.9t - 340) = 0, since t=0 is not going to be it then t=340/4.9 maybe

OpenStudy (amistre64):

somethings not setting right with that to me

OpenStudy (anonymous):

That would lead t to equal 69.39

OpenStudy (amistre64):

yeah ... trying to see what the flaw is if any logically, the distance it takes the rock to fall is the same distance it takes for the sound to travel back, maybe using the same t parameter is a bad idea and we need distance to fall = 4.9 a^2 distance back = 340 b such that a+b = 4.2

OpenStudy (amistre64):

this is a system of equations that may be solvable

OpenStudy (amistre64):

a+b = 4.2; b=4.2-a 4.9a^2 - 340b = 0 4.9a^2 - 340(4.2-a) = 0 now its a quadratic in a

OpenStudy (amistre64):

a = 3.97256 according to the wolf, for speed of a solution that is

OpenStudy (anonymous):

so is that the answer?

OpenStudy (amistre64):

thats what the value of a is, the answer should be d=4.9a^2

OpenStudy (amistre64):

or d = 340(4.2-a)

OpenStudy (anonymous):

d=77.23

OpenStudy (anonymous):

or 78.2

OpenStudy (amistre64):

im getting about 77.33

OpenStudy (anonymous):

ok ill try all 3

OpenStudy (anonymous):

ok i got it!!!! thanks so much for the help

OpenStudy (amistre64):

:) youre welcome, that was kinda fun .... at least it made me think more than usual

OpenStudy (anonymous):

yes it is a very challenging high school course!